# On the Convergence of Tâtonnement for Linear Fisher Markets

### v1

@@ Appendix D Omitted proofs in Sections 4 and 5
By Corollary 1, we have $\lVert g^{t}\rVert\leq G$ for any $g^{t}\in\partial\varphi(p^{t})$ for all $t\geq 0$. By Eq. 22, we have that after $T$ iterations,
$\varphi(p^{T})-\varphi(p^{*})$ $\leq\frac{GR}{\sqrt{T}}\left(\frac{5}{4}+\frac{1}{4}\log{T}\right)$
$\leq\epsilon\frac{\frac{5}{4}+\frac{1}{2}\log(\frac{GR}{\epsilon})+\frac{1}{2}\log{\log(\frac{GR}{\epsilon})}}{\log(\frac{GR}{\epsilon})}$
@@
where the third inequality follows by $\log(\frac{GR}{\epsilon})\geq 4$ and $\frac{\log{x}}{x}\leq\frac{1}{e}$ for all $x>0$.
D.2 Proof of Lemma 5.
@@
Let $\hat{p}$ be an $m$-dimensional price vector where
@@
Assume that we adjust prices in the quasi-linear Fisher market using Eq. 3, with any initial price $p^{0}\geq\hat{p}$ and stepsize $\eta<\frac{1}{2m}\min_{j}\hat{p}_{j}$. Then, $p^{t}_{j}\geq\hat{p}_{j}-2m\eta$ for all $j\in[m]$ and $t\geq 0$.
First, analogous to the proof of Lemma 1, we can show the following two facts. Recall that buyer $\boldsymbol{b}(j)$ is defined as the buyer with the minimum index such that
$\boldsymbol{b}(j)\in\operatorname*{\textnormal{Argmax}}_{i}\frac{v_{ij}}{\|v_{i}\|_{\infty}}.$ (36)
@@
Now we are ready to prove Lemma 5.
@@
We suppose that there exists a $T=\min\big{\{}t\,\big{|}\,p^{t}_{j}<\hat{p}_{j}-2m\eta\text{ for some }j\big{\}}<\infty$, otherwise we are done. In words, $T$ is the first time that some item $j$’s price goes below $\hat{p}_{j}-2m\eta$. For each item $j$ such that $p^{T}_{j}<\hat{p}_{j}-2m\eta$, we let $t_{j}<T$ be the last time that its price went from above $\hat{p}_{j}$ to below $\hat{p}_{j}$. It follows that $p^{t}_{j}\leq\hat{p}_{j}$ for $t_{j}\leq t\leq T$. Since the price decreases by at most $\eta$ per iteration, $T\geq t_{j}+2m$, and thus $p^{t}_{j}\leq\hat{p}_{j}$ for $t_{j}\leq t\leq t_{j}+2m$. Note that, by definition of $T$, $p^{t_{j}}_{j^{\prime}}\geq\hat{p}_{j^{\prime}}-2m\eta$ for all $j^{\prime}\neq j$.
$p^{t_{j}+2m}_{j}\geq p^{t_{j}}_{j}+\eta\sum_{t=t_{j}}^{t_{j}+2m-1}\Big{(}\frac{b^{t}_{\boldsymbol{b}(j)j}}{p^{t}_{j}}-1\Big{)}$ $\geq p^{t_{j}}_{j}+\eta\sum_{t=t_{j}}^{t_{j}+2m-1}\frac{b^{t}_{\boldsymbol{b}(j)j}}{\hat{p}_{j}}-2m\eta$
$\geq p^{t_{j}}_{j}+8m\eta-2m\eta$
$\geq\hat{p}_{j}+(6m-1)\eta,$ (40)
Therefore, Eq. 40 contradicts $p^{t_{j}+2m}_{j}\leq\hat{p}_{j}$ and thus $T<\infty$ does not exist. By contradiction, we can conclude $p^{t}_{j}\geq\hat{p}_{j}-2m\eta$ for all $j$ and all $t\geq 0$. ∎
D.3 Proof of Corollary 3
@@
Let $b_{ij}^{t}$ be the spend from buyer $i$ for item $j$ at time step $t$, then we have
where the last inequality follows because $p^{t}_{j}\geq\min_{j}\hat{p}_{j}-2m\eta$ for all $j$ and $\sum_{j=1}^{m}\sum_{i=1}^{n}b^{t}_{ij}\leq B$. ∎
D.4 Proof of Lemma 6.
Proof.
Then, we consider the largest possible value $p^{t}_{j}$ can attain, approaching from below $\min\{\max_{i}v_{ij},B\}$. For any $p^{t-1}_{j}\leq\min\{\max_{i}v_{ij},B\}$, then $p^{t}_{j}=p^{t-1}_{j}+\eta\left(x_{j}-1\right)\leq p^{t-1}_{j}+\eta\frac{\sum_{i}b^{t}_{ij}}{p^{t-1}_{j}}\leq\min\{\max_{i}v_{ij},B\}+\eta\frac{B}{p^{t-1}_{j}}\leq\min\{\max_{i}v_{ij},B\}+\frac{\eta}{\hat{p}_{j}-2m\eta}B$. Hence, the lemma follows. ∎
D.5 Proof of Theorem 3
Proof.
Since $g^{t}\in\partial\varphi(p^{t})$ is equal to the excess supply $-z^{t}$, by Corollary 3 we have $\lVert g^{t}\rVert=\lVert z^{t}\rVert\leq G:=\frac{B}{\min_{j}\hat{p}_{j}-2m\eta}+m$. By Lemma 7, the quadratic growth parameter associated with $\varphi^{q}(p)$ is $\alpha=\min_{j}\frac{p^{*}_{j}}{2\big{(}\min\{\max_{i}v_{ij},B\}+\frac{\eta B}{\hat{p}_{j}-2m\eta}\big{)}^{2}}$. The rest of the proof follows the same steps as in the proof of Eq. 17. ∎
D.6 Proof of Corollary 4
@@
$\alpha\geq\min_{j}\frac{p^{*}_{j}}{2\big{(}B+\frac{\eta B}{\hat{p}_{j}-2m\eta}\big{)}^{2}}\geq\frac{2\min_{j}p^{*}_{j}}{9B^{2}}.$ (42)
@@
$e=\frac{\eta}{2\alpha}\left(\frac{B}{\min_{j}\hat{p}_{j}-2m\eta}+m\right)^{2}\leq\frac{9\eta B^{2}}{4\min_{j}p^{*}_{j}}\left(\frac{2B}{\min_{j}\hat{p}_{j}}+m\right)^{2}=\frac{\epsilon}{2},$ (43)
@@
where the inequality follows by $\eta\leq\frac{\min_{j^{\prime}}\hat{p}_{j^{\prime}}}{4m}$ and the last equality follows by $\eta\leq\frac{2\epsilon\min_{j}p^{*}_{j}}{9B^{2}(\frac{2B}{\min_{j}\hat{p}_{j}}+m)^{2}}$. After $t=\frac{1}{2\eta\alpha}\log{\frac{8B^{2}}{\epsilon}}$ iterations, we have
$t\log{(1-2\eta\alpha)}+\log{\lVert p^{0}-p^{*}\rVert^{2}}\leq-2\eta\alpha\cdot t+\log{4B^{2}}=-\log{\frac{8B^{2}}{\epsilon}}+\log{4B^{2}}=\log{\frac{\epsilon}{2}},$ (44)
@@
where the inequality follows because $\log(1+x)\leq x,\;\forall\,x>-1$ and $\lVert p^{0}-p^{*}\rVert\leq 2B$. Thus, by Eq. 17 $\lVert p^{t}-p^{*}\rVert^{2}\leq(1-2\eta\alpha)^{t}\lVert p^{0}-p^{*}\rVert^{2}+e\leq\frac{\epsilon}{2}+\frac{\epsilon}{2}=\epsilon$.
@@
Since $\alpha\geq\frac{2\min_{j}p^{*}_{j}}{9B^{2}}$, $t\leq\frac{9B^{2}}{4\eta\min_{j}p^{*}_{j}}\log(\frac{8B^{2}}{\epsilon})$. If $\epsilon$ is small such that $\frac{2\epsilon\min_{j}p^{*}_{j}}{9B^{2}(\frac{2B}{\min_{j}\hat{p}_{j}}+m)^{2}}\leq\frac{\min_{j}\hat{p}_{j}}{4m}$, $\eta$ is of the order $\mathcal{O}(\epsilon)$, hence the time complexity is of the order $\mathcal{O}(\frac{1}{\epsilon}\log{\frac{1}{\epsilon}})$. Otherwise, $\eta$ is independent of $\epsilon$, hence the time complexity is of the order $\mathcal{O}(\log{\frac{1}{\epsilon}})$. ∎

### v2

@@ Appendix C More Detailed Discussion on the LFM
By Lemma 1 $(ii)$, we have $\lVert g^{t}\rVert\leq G$ for any $g^{t}\in\partial\varphi(p^{t})$ for all $t\geq 0$. By Eq. 26, we have that after $T$ iterations,
$\varphi(p^{T})-\varphi(p^{\textup{*}})\leq\frac{GR}{\sqrt{T}}\left(\frac{5}{4}+\frac{1}{4}\log{T}\right)$ $\leq\epsilon\frac{\frac{5}{4}+\frac{1}{2}\log(\frac{GR}{\epsilon})+\frac{1}{2}\log{\log(\frac{GR}{\epsilon})}}{\log(\frac{GR}{\epsilon})}$
@@
where the third inequality follows by $\log(\frac{GR}{\epsilon})\geq 4$ and $\frac{\log{x}}{x}\leq\frac{1}{e}$ for all $x>0$. ∎
Discussion on constant versus adaptive stepsize settings.
In this paper, we focus on the constant stepsize setting because it aligns better with the natural economic dynamics. From the view of economics, the stepsize in tâtonnement corresponds to a “speed of adjustment” parameter, which is a simple characteristic of the market and commonly seen as a constant. See, for example, Bala and Majumdar (1992) (page 5, after Eq. (3.1)) and  Cole and Fleischer (2008) (page 4, after Eq. (2)). If we allow for adaptive stepsizes, then there are results showing that, under a quadratic growth condition, the subgradient descent method will converge to the exact optimal solution (e.g., decaying stepsizes in Johnstone and Moulin (2020), Polyak stepsizes in Grimmer (2019), etc.). As a result, tâtonnement converges to the exact market equilibrium under such adaptive stepsize schemes.
Appendix D Proofs in Section 5
First, we show that the tâtonnement process in the QLFM generates a sequence of prices with lower and upper bounds, as long as the stepsize is not too large. Lower bounds on prices in the QLFM can be obtained almost identically to how we prove Lemma 1 since QLFM behaves the same as LFM when the price is small. Analogous to the LFM case, the upper bounds on the excess demand and prices follow as well. Recall that, given an item $j$, we use $\boldsymbol{b}(j)$ to denote the buyer with the minimum index such that
$\boldsymbol{b}(j)\in\operatorname\textup{*}{\textnormal{Argmax}}_{i}\frac{v_{ij}}{\|v_{i}\|_{\infty}}.$
@@ Appendix D Proofs in Section 5
Let $v_{\min}=\min_{j}v_{\boldsymbol{b}(j)j}$. Let $\hat{p}$ be an $m$-dimensional price vector where
@@
Assume that we adjust prices in the QLFM with Eq. 4, starting from any initial price $p^{0}\geq\hat{p}$, with any stepsize $\eta<\frac{1}{2m}\min_{j}\hat{p}_{j}$. Then, we have for all $t\geq 0$,
  • (i)
    $p^{t}_{j}\geq\hat{p}_{j}-2m\eta$ for all $j\in[m]$;
  • (ii)
    $\lVert z(p^{t})\rVert\leq\frac{B}{\min_{j}\hat{p}_{j}-2m\eta}+m$;
  • (iii)
    $p^{t}_{j}\leq\min\{\max_{i}v_{ij},B\}+\frac{\eta B}{\hat{p}_{j}-2m\eta}$ for all $j\in[m]$.
Proof.
  • (i)
    Recall that, given an item $j$, we use $\boldsymbol{b}(j)$ to denote the buyer with the minimum index such that
    $\boldsymbol{b}(j)\in\operatorname\textup{*}{\textnormal{Argmax}}_{i}\frac{v_{ij}}{\|v_{i}\|_{\infty}}.$
    As in the proof of Lemma 1 $(i)$, we can show two facts analogous to Facts 1 and 17. Moreover, $\hat{p}_{j}<v_{\min}\max_{i}\frac{v_{ij}}{\lVert v_{i}\rVert_{\infty}}\leq v_{\boldsymbol{b}(j)j}$ guarantees buyer $\boldsymbol{b}(j)$ wants to spend all her money if $p_{j}\leq\hat{p}_{j}$. Other parts of the proof are analogous to the proof of Lemma 1 $(i)$. We show a complete proof as follows.
    First, analogous to the proof of Lemma 1, we can show the following two facts.
@@
    Now we are ready to prove Lemma 5 $(i)$. We suppose that there exists a $T=\min\big{\{}t\,\big{|}\,p^{t}_{j}<\hat{p}_{j}-2m\eta\text{ for some }j\big{\}}<\infty$, otherwise we are done. In words, $T$ is the first time that some item $j$’s price goes below $\hat{p}_{j}-2m\eta$. For each item $j$ such that $p^{T}_{j}<\hat{p}_{j}-2m\eta$, we let $t_{j}<T$ be the last time that its price went from above $\hat{p}_{j}$ to below $\hat{p}_{j}$. It follows that $p^{t}_{j}\leq\hat{p}_{j}$ for $t_{j}\leq t\leq T$. Since the price decreases by at most $\eta$ per iteration, $T\geq t_{j}+2m$, and thus $p^{t}_{j}\leq\hat{p}_{j}$ for $t_{j}\leq t\leq t_{j}+2m$. Note that, by definition of $T$, $p^{t_{j}}_{j^{\prime}}\geq\hat{p}_{j^{\prime}}-2m\eta$ for all $j^{\prime}\neq j$.
    $p^{t_{j}+2m}_{j}\geq p^{t_{j}}_{j}+\eta\sum_{t=t_{j}}^{t_{j}+2m-1}\Big{(}\frac{b^{t}_{\boldsymbol{b}(j)j}}{p^{t}_{j}}-1\Big{)}\geq p^{t_{j}}_{j}+\eta\sum_{t=t_{j}}^{t_{j}+2m-1}\frac{b^{t}_{\boldsymbol{b}(j)j}}{\hat{p}_{j}}-2m\eta\geq p^{t_{j}}_{j}+8m\eta-2m\eta$ $\geq\hat{p}_{j}+(6m-1)\eta,$ (34)
    Therefore, Eq. 34 contradicts $p^{t_{j}+2m}_{j}\leq\hat{p}_{j}$ and thus $T<\infty$ does not exist. By contradiction, we can conclude $p^{t}_{j}\geq\hat{p}_{j}-2m\eta$ for all $j$ and all $t\geq 0$.
  • (ii)
    Then, the part $(ii)$ follows.
    where the last inequality follows because $p^{t}_{j}\geq\min_{j}\hat{p}_{j}-2m\eta$ for all $j$ and $\sum_{j=1}^{m}\sum_{i=1}^{n}b^{t}_{ij}\leq B$.
  • (iii)
    Then, we consider the largest possible value $p^{t}_{j}$ can attain, approaching from below $\min\{\max_{i}v_{ij},B\}$. For any $p^{t-1}_{j}\leq\min\{\max_{i}v_{ij},B\}$, then $p^{t}_{j}=p^{t-1}_{j}+\eta\left(x_{j}-1\right)\leq p^{t-1}_{j}+\eta\frac{\sum_{i}b^{t}_{ij}}{p^{t-1}_{j}}\leq\min\{\max_{i}v_{ij},B\}+\eta\frac{B}{p^{t-1}_{j}}\leq\min\{\max_{i}v_{ij},B\}+\frac{\eta}{\hat{p}_{j}-2m\eta}B$. Hence, the part $(iii)$ follows.
Lemma 6.
The convex function $\varphi^{q}(p)$ satisfies the quadratic growth condition with modulus
$\alpha=\min_{j}\frac{p^{\textup{*}}_{j}}{2\big{(}\min\{\max_{i}v_{ij},B\}+\frac{\eta B}{\hat{p}_{j}-2m\eta}\big{)}^{2}},$
where $*p^{\textup{*}}$ denotes the unique equilibrium price vector and $\hat{p}$ is defined in Lemma 5.
@@
Let $*(p^{\textup{*}},x^{\textup{*}},y^{\textup{*}})$ be a market equilibrium in the QLFM and
$h^{q}(p)=\sum_{j=1}^{m}p_{j}-\sum_{i=1}^{n}\sum_{j=1}^{m}p^{\textup{*}}_{j}x^{\textup{*}}_{ij}\log{\left(\frac{p_{j}}{v_{ij}}\right)}.$ (36)
By KKT conditions of the EG-like primal and dual problems for QLFM, we have $\langle p^{\textup{*}},x^{\textup{*}}_{i}\rangle+y^{\textup{*}}_{i}=B_{i}$ and $y^{\textup{*}}_{i}>0$ implies $\min_{k\in[m]}\frac{p_{k}}{v_{ik}}\geq 1$ for all $i\in[n]$. Thereby, we have
$\varphi^{q}(p)$ $=\sum_{j=1}^{m}p_{j}-\sum_{i=1}^{n}\left(\sum_{j=1}^{m}p^{\textup{*}}_{j}x^{\textup{*}}_{ij}+y_{i}^{\textup{*}}\right)\log{\left(\min\left\{\min_{k\in[m]}\frac{p_{k}}{v_{ik}},1\right\}\right)}$
$=\sum_{j=1}^{m}p_{j}-\sum_{i=1}^{n}\sum_{j=1}^{m}p^{\textup{*}}_{j}x^{\textup{*}}_{ij}\log{\left(\min\left\{\min_{k\in[m]}\frac{p_{k}}{v_{ik}},1\right\}\right)}-\sum_{i}y_{i}^{\textup{*}}\log{\left(\min\left\{\min_{k\in[m]}\frac{p_{k}}{v_{ik}},1\right\}\right)}$
$=\sum_{j=1}^{m}p_{j}-\sum_{i=1}^{n}\sum_{j=1}^{m}p^{\textup{*}}_{j}x^{\textup{*}}_{ij}\log{\left(\min\left\{\min_{k\in[m]}\frac{p_{k}}{v_{ik}},1\right\}\right)}\geq h^{q}(p).$ (37)
Also, $*\varphi^{q}(p^{\textup{*}})=h^{q}(p^{\textup{*}})$ since $x^{\textup{*}}_{ij}>0$ implies $\frac{p^{\textup{*}}_{j}}{v_{ij}}=\min_{k\in[m]}\frac{p^{\textup{*}}_{k}}{v_{ik}}\leq 1$. Analogous to the LFM case, $\nabla_{p_{j}}h^{q}(p)=1-\frac{p^{\textup{*}}_{j}}{p_{j}}\sum_{i=1}^{n}x^{\textup{*}}_{ij}=1-\frac{p^{\textup{*}}_{j}}{p_{j}}$ for all $j\in[m]$, thus $*p=p^{\textup{*}}$ is the minimizer of $h^{q}(p)$. Therefore, $\varphi^{q}(p)-\varphi^{q}(p^{\textup{*}})\geq h^{q}(p)-h^{q}(p^{\textup{*}})\geq\alpha\lVert p-p^{\textup{*}}\rVert^{2}$, where $\alpha=\min_{j}\frac{p^{\textup{*}}_{j}}{2\bar{p}_{j}^{2}}$. ∎
Theorem 2.
$\lVert p^{t}-p^{\textup{*}}\rVert^{2}\leq(1-2\eta\alpha)^{t}\lVert p^{0}-p^{\textup{*}}\rVert^{2}+e\quad\text{for all }t\geq 0,$ (38)
where $\alpha$ is defined in Lemma 6 and
@@
Since $g^{t}\in\partial\varphi(p^{t})$ is equal to the excess supply $-z^{t}$, by Lemma 5 $(ii)$, we have $\lVert g^{t}\rVert=\lVert z^{t}\rVert\leq G:=\frac{B}{\min_{j}\hat{p}_{j}-2m\eta}+m$. By Lemma 6, the quadratic growth parameter associated with $\varphi^{q}(p)$ is $\alpha=\min_{j}\frac{p^{\textup{*}}_{j}}{2\big{(}\min\{\max_{i}v_{ij},B\}+\frac{\eta B}{\hat{p}_{j}-2m\eta}\big{)}^{2}}$. The rest of the proof follows the same steps as in the proof of Theorem 1. ∎
Corollary 2.
For a given $\epsilon>0$, let $\eta\leq\min\{\frac{\min_{j}\hat{p}_{j}}{4m},\frac{2\epsilon\min_{j}p^{\textup{*}}_{j}}{9B^{2}(\frac{2B}{\min_{j}\hat{p}_{j}}+m)^{2}}\}$. Then, starting from any $p^{0}\geq\hat{p}$ such that $\sum_{j}p^{0}_{j}=B$, tâtonnement with stepsize $\eta$ generates a price vector $p$ such that $\lVert p-p^{\textup{*}}\rVert^{2}\leq\epsilon$ in $\mathcal{O}(\frac{1}{\epsilon}\log{\frac{1}{\epsilon}})$ iterations. If $\epsilon\geq\frac{9B^{2}\min_{j}\hat{p}_{j}}{8m\min_{j}p^{\textup{*}}_{j}}(\frac{2B}{\min_{j}\hat{p}_{j}}+m)^{2}$, the time complexity is on the order of $\mathcal{O}(\log{\frac{1}{\epsilon}})$.
@@
$\alpha\geq\min_{j}\frac{p^{\textup{*}}_{j}}{2\big{(}B+\frac{\eta B}{\hat{p}_{j}-2m\eta}\big{)}^{2}}\geq\frac{2\min_{j}p^{\textup{*}}_{j}}{9B^{2}}.$ (39)
@@
$e=\frac{\eta}{2\alpha}\left(\frac{B}{\min_{j}\hat{p}_{j}-2m\eta}+m\right)^{2}\leq\frac{9\eta B^{2}}{4\min_{j}p^{\textup{*}}_{j}}\left(\frac{2B}{\min_{j}\hat{p}_{j}}+m\right)^{2}=\frac{\epsilon}{2},$ (40)
@@
where the inequality follows by $\eta\leq\frac{\min_{j^{\prime}}\hat{p}_{j^{\prime}}}{4m}$ and the last equality follows by $\eta\leq\frac{2\epsilon\min_{j}p^{\textup{*}}_{j}}{9B^{2}(\frac{2B}{\min_{j}\hat{p}_{j}}+m)^{2}}$. After $t=\frac{1}{2\eta\alpha}\log{\frac{8B^{2}}{\epsilon}}$ iterations, we have
$t\log{(1-2\eta\alpha)}+\log{\lVert p^{0}-p^{\textup{*}}\rVert^{2}}\leq-2\eta\alpha\cdot t+\log{4B^{2}}=-\log{\frac{8B^{2}}{\epsilon}}+\log{4B^{2}}=\log{\frac{\epsilon}{2}},$ (41)
@@
where the inequality follows because $\log(1+x)\leq x,\;\forall\,x>-1$ and $\lVert p^{0}-p^{\textup{*}}\rVert\leq 2B$. Thus, by Theorem 1 $\lVert p^{t}-p^{\textup{*}}\rVert^{2}\leq(1-2\eta\alpha)^{t}\lVert p^{0}-p^{\textup{*}}\rVert^{2}+e\leq\frac{\epsilon}{2}+\frac{\epsilon}{2}=\epsilon$.
@@
Since $\alpha\geq\frac{2\min_{j}p^{\textup{*}}_{j}}{9B^{2}}$, $t\leq\frac{9B^{2}}{4\eta\min_{j}p^{\textup{*}}_{j}}\log(\frac{8B^{2}}{\epsilon})$. If $\epsilon$ is small such that $\frac{2\epsilon\min_{j}p^{\textup{*}}_{j}}{9B^{2}(\frac{2B}{\min_{j}\hat{p}_{j}}+m)^{2}}\leq\frac{\min_{j}\hat{p}_{j}}{4m}$, $\eta$ is of the order $\mathcal{O}(\epsilon)$, hence the time complexity is of the order $\mathcal{O}(\frac{1}{\epsilon}\log{\frac{1}{\epsilon}})$. Otherwise, $\eta$ is independent of $\epsilon$, hence the time complexity is of the order $\mathcal{O}(\log{\frac{1}{\epsilon}})$. ∎
Appendix E Numerical Experiments Details and More Results
Real data description.
For the real data experiments, we implement tâtonnement on a market instance used in Nan, Gao, and Kroer (2023). Different from their unit budgets, we consider randomly generated budgets from the uniform distribution on $[0,1]$. For the linear Fisher market, we normalize the budgets such that the total budget is $1$; for the quasi-linear Fisher market, we use the non-normalized budgets, as money has value outside the current market under QL setting. We assign each item a unit supply.
The valutions in the market instance is generated by using a movie rating dataset collected from twitter called Movietweetings (Dooms, De Pessemier, and Martens 2013) (“snapshots 200K”). Here, users are viewed as the buyers, movies as items, and ratings as valuations. As in Nan, Gao, and Kroer (2023), since the data in the original dataset is sparse, we remove users and movies with too few entries. We then complete the matrix by using the matrix completion software fancyimpute (Rubinsteyn and Feldman 2016). We normalize the utilities such that the sum of utilities of each buyer is $1$. Finally, the resulting instance has $n=691$ buyers and $m=632$ items. We use random seed $1$ to generate budgets for buyers.
In LFM setting, we test the convergence of tâtonnement with three stepsizes: $\eta=1\times 10^{-6},2\times 10^{-6},3\times 10^{-6}$; in QLFM setting, we test the convergence with three stepsizes: $\eta=1\times 10^{-7},2\times 10^{-7},3\times 10^{-7}$. We run $8000$ and $70000$ iterations for LFM and QLFM, respectively, to guarantee tâtonnement converges to a small neighborhood of the equilibrium price.
More detailed description of synthetic data experiments.
In the synthetic data experiments, we consider five different distributions of utilities:
  • Uniform: $v$ is generated from the uniform distribution on $[0,1]$.
  • Log-normal: $v$ is generated from the log-normal distribution associated with the standard normal distribution $\mathcal{N}(0,1)$.
  • Exponential: $v$ is generated from the exponential distribution with the scale parameter $1$.
  • Truncated normal: $v$ is generated from the truncated normal distribution associated with $\mathcal{N}(0,1)$, and truncated at ${10}^{-3}$ and $10$ standard deviations.
  • Uniform integers: $v$ is generated from the uniform distribution on $\{1,\ldots,100\}$.
For each distribution, we generate four instances of different sizes: $(n,m)=(10,20),(20,40),(30,60),(40,80)$, where $n$ is the number of buyers and $m$ is the number of items. We use random seed $1$ to generate all budgets and valuation matrices. For all synthetic instances, we test the convergence of tâtonnement with three stepsizes: $\eta=2\times 10^{-5},4\times 10^{-5},6\times 10^{-5}$. We run $200000$ iterations for each instance to guarantee tâtonnement converges to a small neighborhood of the equilibrium price. Note that we only show a part of iterations in the plots for better visualization. To more clearly show linear convergence, we zoom in on the “straight line” portion of the plots (the initial part of all iterations) in each figure.
More details on the experiments.
In our experiments, we choose the MBB item with the smallest index when there is more than one MBB item. In all instances, we run tâtonnement starting with $p^{0}_{j}=\frac{1}{m}$ for all $j\in[m]$, where $m$ is the number of items. We use the squared error norm to measure the convergence because it is consistent with our theoretical results. All experiments are conducted on a personal computer using Python 3.9.12. For our hyperparameter settings, we tried around $10$ stepsizes ($1\times 10^{-7}-1\times 10^{-4}$) for each setting and chose those that can clearly show the convergence.
Figure 4: Convergence of squared error norms on random generated instances ($v$ is generated from the uniform distribution $[0,1)$) of different sizes under linear utilities.
Figure 5: Convergence of squared error norms on random generated instances ($v$ is generated from the uniform distribution $[0,1)$) of different sizes under quasi-linear utilities.
Figure 6: Convergence of squared error norms on random generated instances ($v$ is generated from the log-normal distribution associated with the standard normal distribution $\mathcal{N}(0,1)$) of different sizes under linear utilities.
Figure 7: Convergence of squared error norms on random generated instances ($v$ is generated from the log-normal distribution associated with the standard normal distribution $\mathcal{N}(0,1)$) of different sizes under quasi-linear utilities.
Figure 8: Convergence of squared error norms on random generated instances ($v$ is generated from the exponential distribution with the scale parameter $1$) of different sizes under linear utilities.
Figure 9: Convergence of squared error norms on random generated instances ($v$ is generated from the exponential distribution with the scale parameter $1$) of different sizes under quasi-linear utilities.
Figure 10: Convergence of squared error norms on random generated instances ($v$ is generated from the truncated normal distribution associated with $\mathcal{N}(0,1)$, and truncated at ${10}^{-3}$ and $10$ standard deviations from 00) of different sizes under linear utilities.
Figure 11: Convergence of squared error norms on random generated instances ($v$ is generated from the truncated normal distribution associated with $\mathcal{N}(0,1)$, and truncated at ${10}^{-3}$ and $10$ standard deviations from 00) of different sizes under quasi-linear utilities.
Figure 12: Convergence of squared error norms on random generated instances ($v$ is generated from the uniform random integers on $\{1,\ldots,100\}$) of different sizes under linear utilities.
Figure 13: Convergence of squared error norms on random generated instances ($v$ is generated from the uniform random integers on $\{1,\ldots,100\}$) of different sizes under quasi-linear utilities.
Appendix F Comparison of different types of Tâtonnement.
We compare the convergence of the coordinate tâtonnement and the standard tâtonnement on the real data instance. See Figs. 14, 15, 16, 17 and 18 for the results on synthetic indtances and see Fig. 19 for the results on the real data instance.
Experimental details.
We mainly compare the convergence of
  • Additive tâtonnement: $p^{t+1}_{j}=p^{t}_{j}+\eta z(p^{t})$;
  • Multiplicative tâtonnement: $p^{t+1}_{j}=p^{t}_{j}(1+\eta z(p^{t}))$;
  • Entropic tâtonnement: $p^{t+1}_{j}=p^{t}_{j}\exp(\eta z(p^{t}))$.
Here, $\eta$ is the stepsize, and $z(p^{t})$ is the excess demand vector at time $t$. To be fair, we use a constant stepsize $\eta$ for additive tâtonnement and $\eta^{\prime}=\frac{m\eta}{\lVert B\rVert_{1}}$ for the multiplicative and entropic tâtonnement to ensure the stepsize is of the similar scale as the additive tâtonnement. In particular, we use $\eta=2\times 10^{-5}$ for random generated instances and $\eta=2\times 10^{-6}$ for the real data instance for the linear Fisher market; we use $\eta=2\times 10^{-5}$ for random generated instances and $\eta=2\times 10^{-7}$ for the real data instance for the quasi-linear Fisher market.
Discussion.
As shown in the figures, the three types of tâtonnement have similar convergence behaviors: they all converge to the equilibrium price in a linear rate, and then oscillate around the equilibrium price. The multiplicative and entropic tâtonnement are almost the same, which is as expected since multiplicative tâtonnement is an approximation of entropic tâtonnement when the stepsize is small and the excess demand is bounded. In the linear Fisher market, the additive tâtonnement has a slightly slower convergence rate than the other two in some instances; in the quasi-linear Fisher market, the additive tâtonnement has a faster convergence rate than the other two in most instances. This is an interesting observation, and we leave the more structural comparison of different types of tâtonnement for future work. Note that the convergence speed of tâtonnements might change as we tune the stepsize.
More variants of Tâtonnement.
As shown in Table 1, there are some modified versions of the above three types of tâtonnement in the literature. There are two common modifications:
  • (1)
    Upper bounded the excess demand vector: $z(p^{t})=\min\{z(p^{t}),1\}$;
  • (2)
    Lower bounded the price vector with a small value: $p^{t+1}_{j}=\max\{p^{t}_{j}+\eta z(p^{t}),\epsilon\}$, where $\epsilon$ is a small value.
Via our experiments, we find that the convergence of tâtonnement with $(2)$ is similar to the standard tâtonnement. However, tâtonnement with $(1)$ usually lead to a larger final error norm.
Figure 14: Comparison of different types of Tâtonnement on random generated instances ($v$ is generated from the uniform distribution $[0,1)$) of different sizes under linear utilities (upper row) and quasi-linear utilities (lower row).
Figure 15: Comparison of different types of Tâtonnement on random generated instances ($v$ is generated from the log-normal distribution associated with the standard normal distribution $\mathcal{N}(0,1)$) of different sizes under linear utilities (upper row) and quasi-linear utilities (lower row).
Figure 16: Comparison of different types of Tâtonnement on random generated instances ($v$ is generated from the exponential distribution with the scale parameter $1$) of different sizes under linear utilities (upper row) and quasi-linear utilities (lower row).
Figure 17: Comparison of different types of Tâtonnement on random generated instances ($v$ is generated from the truncated normal distribution associated with $\mathcal{N}(0,1)$, and truncated at ${10}^{-3}$ and $10$ standard deviations from 00) of different sizes under linear utilities (upper row) and quasi-linear utilities (lower row).
Figure 18: Comparison of different types of Tâtonnement on random generated instances ($v$ is generated from the uniform random integers on $\{1,\ldots,100\}$) of different sizes under linear utilities (upper row) and quasi-linear utilities (lower row).
Figure 19: Comparison of different types of Tâtonnement on the real data instance under linear utilities (left) and quasi-linear utilities (right).